# Knowledge Distillation for Underwater Feature Extraction and Matching via GAN-synthesized Images

### v1

@@ V-B GAN Synthetic Results
V-B1 Naive Case - Pure Underwater Background Synthesis
To evaluate the performance of the GAN synthesis step, we first conduct a simple qualitative experiment where the target underwater images contain only the pure underwater background. In this scenario, $J_{c}$ in (2) is set to 0, simplifying the formation model in (7) to:
$I^{Pure}_{c}(x)=B_{c}^{\infty}+B_{c}^{Noise}=B_{c}^{\infty}(1-M(x))+M(x),$ (19)
we extracted two distinct underwater environments from the marine snow dataset released in [19], which consists of footage with pure backgrounds, to serve as the target image $T_{c}$. The visual results from the GAN are displayed in Fig. 4. These findings confirm that our method produces a synthetic noise distribution akin to the marine snow in the target environment and accurately estimates the underwater background lights $B_{c}^{\infty}$. Therefore, our approach effectively estimates the underwater background lights and the noise distribution within the target environment.
V-B2 Target Underwater Image Synthesis
To evaluate the complete synthesis of target environments, we selected two underwater sequences: the EASI-T sequence, sourced from the EASI dataset [11], which was created in an indoor water tank with baby powder used to adjust turbidity; and the Lake-C sequence, a simulated cemetery scene captured in a lake, sourced from [14]. Qualitative examples are shown in Fig. 5. We visually demonstrate the difference between the smooth synthetic images without GAN-synthesized noise and forward scattering and the fully synthetic images. To assess the effectiveness of each component in the synthesis process, we conducted a quantitative evaluation using the GAN metric Frechet Inception Distance (FID), where a lower score indicates a closer resemblance to the target images. We compared the FID scores of four groups against real underwater images: in-air RGB images ($J_{c}$), smooth synthetic images without noise and forward scattering (Smooth), synthetic images without GAN noise ($I_{c}$), and fully synthesized images ($I^{Noise}_{c}$). Each group consisted of 500 randomly sampled images, with the results presented in Fig. 6. We observe a decreasing trend in the FID scores as components are added, as illustrated in Fig. 6(a). Using the EASI-T sequence as an example, we employed t-SNE to visualize the domain shift, which clearly shows a progression from left to right towards the target images in Fig. 6(b). Notably, adding forward scattering and image noise components moves the synthetic images closer to the target domain.
@@ V-B2 Target Underwater Image Synthesis
(a) Image Pairs
(b) SuperPoint [6]
(c) w/o F. Scattering
(d) Full Synthesized
Figure 7: Qualitative Ablation Results on Feature Matching: Correct matches (yellow), incorrect (blue). With the addition of GAN-synthesized noise and forward scattering, the number of incorrect matches is reduced.
Models Lake-C EASI-T
@@
ALIKE [34] 258.0 0.877 172.4 0.724
ALIKE (KD) 264.0 0.929 220.5 0.844
SuperPoint [6] 153.3 0.592 156.9 0.653
SuperPoint (KD) 275.2 0.904 228.2 0.852
UFEN [11] 244.2 0.807 184.6 0.692
SuperPoint (KD-B) 246.6 0.861 205.1 0.821
TABLE I: Quantitative Evaluation on the Lake-C and EASI-T Sequences: Both ALIKE(KD) and SuperPoint(KD) exhibit improved performance on their respective sequences. The binary model, SuperPoint(KD-B), outperforms UFEN.
V-C Feature Extraction and Matching Results
Using synthetic underwater images as the medium for knowledge distillation, the transferred model demonstrates enhanced performance in feature extraction on images from the target environment. To showcase the generalization of our training framework, we employ not only the SuperPoint [6] as the teacher model as in [11] but also the ALIKE model from [34] as the teacher model. We use the models to extract $500$ feature points from $480\times 640$ images. Due to the lack of ground truth in real underwater sequences, we apply RANSAC to estimate the transformation between image frames with 5-image intervals and filter out mismatches. The performance of the models is evaluated using the Matching Number and the Matching Score metrics. Specifically, Matching Num (M. Num) is defined as $N_{match}$, which denotes the number of correct matches after RANSAC filtering, and Matching Rate (M. Rate) is given by $N_{match}/N_{found}$, where $N_{found}$ represents the total number of matches identified using the descriptors. We randomly sample $100$ image pairs from the sequences EASI-T and Lake-C and get the average, the evaluation results are shown in Table I. We use the ALIKE(KD) and SuperPoint(KD) to represent the models trained by our proposed knowledge distillation framework on the corresponding sequences. Additionally, to demonstrate the effectiveness of the proposed model and its performance compared to UFEN in [11], which was trained with randomly sampled water parameters and image white noise, we also performed knowledge distillation on SuperPoint with a binarization layer and the straight-through gradient estimator, labelled as SuperPoint (KD-B), for a fair comparison with UFEN, whose descriptor is in binary format. The results demonstrate the effectiveness of our proposed knowledge distillation method. The general framework is versatile, working with different in-air feature extraction models and supporting various descriptor formats.
Synthetic Components Performance Metrics
G. Noise GAN Noise F. Scattering M. Num M. Rate
131.1 0.569
166.9 0.727
175.3 0.755
193.2 0.756
228.2 0.852
TABLE II: Ablation study on the effect of forward scattering and noise map: The best performance is achieved when both GAN noise and forward scattering are applied.
Seq. ORB-SLAM3 UFEN-SLAM Ours
Lake-C 11.9 82.1 100
EASI-T 0.0 9.7 100
TABLE III: VSLAM Tracking Proportions (%): UFEN-SLAM shows more robust performance on both sequences.
To evaluate the effectiveness of incorporating forward scattering and image noise in our synthetic image model, we conducted an ablation study using the knowledge-distilled SuperPoint model on the EASI-T sequence (Table II). Both Gaussian image noise (G. Noise) and GAN-synthesized noise (GAN Noise) enhance feature extraction and matching performance, with GAN Noise yielding superior results. Incorporating forward scattering (F. Scattering) further boosts performance. When both GAN-synthesized noise and forward scattering are considered, the model achieves the best performance. This study highlights the importance of accounting for forward scattering and image noise in underwater feature extraction and matching. Qualitative results in Fig. 7 further illustrate the effects of GAN-synthesized noise and forward scattering in image synthesis. Matching correspondences are extracted using pure descriptor matching without RANSAC filtering. We compare the matching performance between the original SuperPoint model and the transferred model trained with synthetic images: one using only GAN noise (w/o F. Scattering) and another using fully synthesized images (Full Synthesized) with both components, where the proportion of incorrect matches is significantly reduced, especially in blurry and background regions.
@@ V-C Feature Extraction and Matching Results
V-D Downstream Application: Underwater VSLAM Results
We further utilized the transferred feature extraction networks, trained using our method, in downstream VSLAM tasks. The knowledge distillation of the SuperPoint pre-trained model was performed using our GAN-synthesized images as the medium, with a binarization layer attached at the end of the descriptor. The model SuperPoint KD-B mentioned above, transferred through the GAN-synthesized images, was integrated into the front end of the ORB-SLAM3 framework, hereafter referred to as Ours, with descriptors converted into a 256-bit binary format to comply with the ORB-SLAM3 standard. The VSLAM performance on the real underwater sequence Lake-C is presented in Fig. 8. Due to the absence of groundtruth for highly turbid underwater sequences, we utilized the Structure-from-Motion method, COLMAP, to reconstruct the Lake-C sequence. However, there are reconstructed frame gaps due to poor matching in blurry regions, as illustrated by image samples from scenes A and B, highlighting substantial reconstruction errors. Both ORB-SLAM3 and UFEN-SLAM failed to track the entire trajectory due to high turbidity and flowing particles, whereas our method maintained tracking throughout the sequence. Similar results were observed on the EASI-T dataset, where even ORB-SLAM3 was unable to initialize. The tracking proportions of these methods on the two sequences are summarized in Tab. III.
@@ VI Conclusion
In this paper, we propose a knowledge distillation framework that transfers in-air feature extraction networks to underwater settings using adaptive GAN-synthesized images. Training on images that account for the forward scattering and image noise, our model outperforms ORB-SALM3 and UFEN-SLAM on the VSLAM task.

### v2

@@ V-B GAN Synthetic Results
V-B1 Target Underwater Image Synthesis
To evaluate the complete synthesis of target environments, we selected the underwater sequences: the EASI-T sequence, sourced from the EASI dataset [11], which was created in an indoor water tank with baby powder used to adjust turbidity and the Lake-C sequence, a simulated cemetery scene captured in a lake, sourced from [14], and the AQUALOC sequence from [47]. Qualitative examples are shown in Fig. 4. We visually demonstrate the difference between the smooth synthetic images without GAN-synthesized noise and forward scattering and the fully synthetic images. To assess the effectiveness of each component in the synthesis process, we conducted a quantitative evaluation using the GAN metric Frechet Inception Distance (FID), where a lower score indicates a closer resemblance to the target images. We compared the FID scores of four groups against real underwater images: in-air RGB images ($J_{c}$), smooth synthetic images without noise and forward scattering (Smooth), synthetic images without GAN noise ($I_{c}$), and fully synthesized images ($I^{Noise}_{c}$). Each group consisted of 500 randomly sampled images, with the results presented in Fig. 5. We observe a decreasing trend in the FID scores as components are added, as illustrated in Fig. 5a. Using the EASI-T sequence as an example, we employed t-SNE to visualize the domain shift, which clearly shows a progression from left to right towards the target images in Fig. 5b. Notably, adding forward scattering and image noise components moves the synthetic images closer to the target domain. While a domain gap remains due to scene content differences between the in-air and underwater datasets, the t-SNE progression, particularly after incorporating image noise, clearly illustrates that modelling noise is both necessary and effective for image synthesis.
TABLE II: Ablation study on the effect of forward scattering and noise
Synthetic Components Performance Metrics
G. Noise GAN Noise F. Scattering M. Num M. Rate
126.9 0.579
104.4 0.468
135.6 0.641
147.5 0.682
152.3 0.642
163.1 0.765
191.3 0.780
@@ V-B1 Target Underwater Image Synthesis
V-C Feature Extraction and Matching Results
Using synthetic underwater images as the medium for knowledge distillation, the transferred model demonstrates enhanced performance in feature extraction and matching on images from the target environment. To showcase the generalization of our training framework, we employ not only the SuperPoint [6] as the teacher model, as in [11] but also the ALIKE model from [36] as the teacher model. We use the models to extract $500$ feature points from $480\times 640$ images. Due to the lack of groundtruth in real underwater sequences, we apply RANSAC to estimate the transformation between image frames with 5-image intervals and filter out mismatches. We evaluate performance using Matching Num (M. N), defined as the number of correct matches $N_{\text{match}}$ after RANSAC filtering, and Matching Rate (M. R), calculated as $N_{\text{match}}/N_{\text{found}}$, where $N_{\text{found}}$ denotes the total number of matches identified by the descriptors. The results are averaged over 100 randomly sampled image pairs from the sequences, as shown in Table I. We use ALIKE(KD) and SP(KD) to represent the models trained by our proposed knowledge distillation framework on the corresponding sequences. Additionally, to demonstrate the effectiveness of the proposed model and its performance compared to UFEN, which was trained with randomly sampled water parameters and image white noise, we also performed knowledge distillation on SuperPoint with a binarization layer, labelled as SP(KD-B), for a fair comparison with UFEN, whose descriptor is in binary format. The results demonstrate the effectiveness of the proposed knowledge distillation framework, which supports various in-air models and descriptor formats.
To evaluate the effectiveness of incorporating forward scattering and image noise in our synthetic image model, we conducted an ablation study using the knowledge-distilled SuperPoint model on the EASI-T sequence (Table II). Both Gaussian image noise (G. Noise) and GAN-synthesized noise (GAN Noise) enhance feature extraction and matching performance, with GAN Noise yielding superior results. Incorporating forward scattering (F. Scattering) further boosts performance. When both GAN-synthesized noise and forward scattering are considered, the model achieves the best performance. This study highlights the importance of accounting for forward scattering and image noise in underwater feature extraction and matching.
TABLE III: Comparison of underwater image synthesis methods
Synthesis Method Lake-C EASI-T
@@ V-C Feature Extraction and Matching Results
WaterGAN 251.9 0.816 140.8 0.649
UWNR 198.5 0.707 104.4 0.457
Ours 275.2 0.904 191.3 0.780
Additionally, to demonstrate the effectiveness of our synthetic method, we compare our GAN-based synthesis with two existing approaches: WaterGAN [22] and UWNR [48] in Tab.III. WaterGAN simulates backscattering generated from random input noise and introduces image noise during generation, but does not model forward scattering. While it brings performance gains for knowledge distillation, our method achieves better results. UWNR simulates the visual appearance of underwater scenes using natural light field rendering, producing visually realistic images. However, it does not account for either image noise or forward scattering, resulting in limited performance in feature transfer. These comparisons highlight the effectiveness of explicitly modeling both image noise and forward scattering in the synthesis pipeline.
V-D Downstream Application: Underwater VSLAM Results
We further utilized the transferred feature extraction networks, trained using our method, in downstream VSLAM tasks. The knowledge distillation of the SuperPoint pre-trained model was performed using our GAN-synthesized images as the medium, with a binarization layer attached at the end of the descriptor. The model SP KD-B mentioned above, transferred through the GAN-synthesized images, was integrated into the front end of the ORB-SLAM3 framework, hereafter referred to as Ours, with descriptors converted into a 256-bit binary format to comply with the ORB-SLAM3 standard. The VSLAM performance on the real underwater sequence Lake-C is presented in Fig. 6. Due to the absence of groundtruth for highly turbid underwater sequences, we utilized the Structure-from-Motion method, COLMAP, to reconstruct the Lake-C sequence. However, there are reconstructed frame gaps due to poor matching in blurry regions, as illustrated by image samples from scenes A and B, highlighting substantial reconstruction errors. Both ORB-SLAM3 and UFEN-SLAM failed to track the entire trajectory due to high turbidity and flowing particles, whereas our method maintained tracking throughout the sequence. Similar results were observed on the EASI-T dataset, where even ORB-SLAM3 was unable to initialize.
TABLE IV: ATE (m) comparison with different models
Model 01 02 03 05 06
SP (w/o KD-B) 0.139 0.0260
UFEN 0.143 0.153 0.0272 0.124 0.0217
Ours 0.118 0.157 0.0259 0.111 0.0211
@@ V-D Downstream Application: Underwater VSLAM Results
Meanwhile, we conducted VSLAM experiments on the AQUALOC dataset to evaluate the transferred model quantitatively. Sequences 04 and 07 were excluded due to consistent tracking failures across all methods, caused by severe lens occlusion and motion-induced blur. To directly reflect the performance of front-end feature detection and matching, we disabled both loop-closure detection and motion mode tracking in the SLAM framework. Feature correspondences were established purely based on descriptor similarity using nearest neighbour matching. We compared the Absolute Trajectory Error (ATE) obtained from VSLAM using three models, a fine-tuned SuperPoint model with descriptor binarization but without knowledge distillation (SP w/o KD-B), UFEN, and our proposed method (SP KD-B). The results are summarized in Tab. IV, indicating that the SuperPoint model without knowledge distillation exhibits unstable tracking performance and larger errors. In contrast, the transferred model trained with our proposed synthesis method, which incorporates both marine snow noise and forward scattering, demonstrates superior performance.
@@ VI Conclusion
In this paper, to improve the feature matching performance of the in-air models in underwater environments, we propose a knowledge distillation framework that adapts in-air feature extraction models to underwater environments using GAN-synthesized images. However, a potential limitation of our method lies in extremely low-texture underwater environments, where the absence of visual cues may hinder matching robustness. Future work may explore integrating inertial or acoustic sensing to enhance performance under such conditions.